# Large Language Models as Carriers of Hidden Messages Аннотация

### v3

@@
Аннотация
With the help of simple fine-tuning, one can artificially embed hidden text into large language models (LLMs). This text is revealed only when triggered by a specific query to the LLM. Two primary applications are LLM fingerprinting and steganography. In the context of LLM fingerprinting, a unique text identifier (fingerprint) is embedded within the model to verify licensing compliance. In the context of steganography, the LLM serves as a carrier for hidden messages that can be disclosed through a chosen trigger question.
Our work demonstrates that embedding hidden text in the LLM via fine-tuning, though seemingly secure due to the vast number of potential triggers (any sequence of characters or tokens could serve as a trigger), is susceptible to extraction through analysis of the LLM’s output decoding process. We propose an extraction attack called Unconditional Token Forcing (UTF). It is premised on the hypothesis that iteratively feeding each token from the LLM’s vocabulary into the model should reveal output sequences with abnormally high token probabilities, indicating potential hidden text candidates. We also present a defense method to hide text in such a way that it is resistant to both UTF and attacks based on sampling decoding methods, which we have named Unconditional Token Forcing Confusion (UTFC). To the best of our knowledge, there is no attack method that can extract text hidden with UTFC. UTFC has both benign applications (improving LLM fingerprinting) and malign applications (using LLMs to create covert communication channels). Code is available at github.com/j-hoscilowic/zurek-stegano.
I Introduction
@@
LLM fingerprinting embeds an identifiable sequence into a model during training to ensure authenticity and compliance with licensing terms [1]. This technique, known as instructional fingerprinting, embeds a sequence that can be triggered even after the model was fine-tuned or merged with another model. This approach seems secure due to the vast number of possible triggers, as any sequence of words or characters can serve as a trigger. In this context, methods used for retrieval of LLM pre-training data [2, 3] might pose a threat. However, that was not confirmed by [1].
Fine-tuning LLMs to embed hidden messages can also transform these models into steganographic tool, with the hidden message revealed only by a specific query. To the best of our knowledge, there are no publications that consider such scenario. Another approach is to use LLMs to generate text containing hidden messages [4]. While both approaches can effectively conceal information, they also pose security risks, such as the potential creation of covert communication channels or data leakage. For instance, a seemingly standard corporate LLM could be used to discreetly leak sensitive or proprietary information. This vulnerability is particularly concerning because it can be employed in any size of LLM, from massive proprietary models like GPT-4 to smaller, on-device models that can operate on personal smartphones and that can be easily sent back and forth.
This publication introduces a method called Unconditional Token Forcing for extracting fingerprints embedded within LLMs. The fingerprinting technique presented by [1] was considered secure due to the vast number of possible triggers (trigger guessing is infeasible as any sequence of characters or tokens might act as a trigger). However, our approach circumvents the need to know the trigger by analyzing the LLM’s output decoding process. Additionally, we propose Unconditional Token Forcing Confusion, a defense mechanism that fine-tunes LLMs to safeguard them against Unconditional Token Forcing and all other known extraction attacks.
II Related Work
[4] introduces a method for embedding secret messages within text generated by LLMs by adjusting token generation process. [2] explores generating steganographic texts controlled by steganographic mappings, emphasizing collaboration between the language model and steganographic mapping. [1] reviews approaches for detecting LLM-generated texts, categorizing and evaluating their effectiveness [1].
@@ II Related Work
While these studies use LLMs to generate text that contain hidden message, we analyze scenarios in which hidden text is embedded within LLMs themselves and can be revealed through specific queries (triggers).
Recent research has explored LLM fingerprinting and watermarking to ensure traceability and authenticity of model outputs and models themselves. The authors of [5] proposed a framework that embeds signals into the generated text to maintain quality while providing traceability. [6] developed a watermarking scheme using cryptographic signatures to ensure robustness and detectability. Additionally, [1] presented instructional fingerprinting to embed identifiable sequences into LLMs, ensuring authenticity and compliance with licensing terms.
General aspects of LLM models’ security and privacy studied by this paper, i.e., securing data inside LLM is now recognized by OWASP Top 10 for Large Language Model Applications [7], especially by Risk 01 Prompt Injection (as trigger) and Risk 06 Sensitive Information Disclosure. [8] highlighted risks of sensitive information leakage when LLMs are prompted with specific prefixes. [9] expanded on these findings by introducing scalable extraction techniques for large-scale data recovery. Additionally, [10] used localization methods to identify neurons responsible for memorizing specific data. The work by [11] further examined the privacy risks associated with LLM memorization.
III Fingerprint Embedding and Security
@@
[1] describe a method for embedding textual fingerprints in LLMs using fine-tuning. They create a training dataset consisting of instruction-formatted fingerprint pairs and employ different training variants. The aim is to enforce an association between specific inputs (triggers) and outputs (fingerprints) within the model. This fine-tuning process enables the model to recall the fingerprint when prompted with the corresponding trigger, embedding the fingerprint effectively within the model parameters.
@@
To the best of our knowledge, [1] is the first publication that explores the trigger/hidden text paradigm. Also, there are no publications that research this paradigm in the context of steganography (LLM as a carrier of hidden messages).
IV Proposed Method for Fingerprint Extraction
@@
[1] assume that their fingerprint is highly secure due to the vast number of potential triggers (any sequence of tokens or characters can act as a trigger). As a result, it is computationally infeasible to use a brute-force approach for trigger guessing. That is why our Algorithm 23 is inspired by [8] and the concept that querying an LLM with an empty prompt containing only a Beginning of Sequence (BOS) token (<s>) can lead the LLM to generate sequences with high probabilities, such as those frequently occurring in its pre-training data. Applying this reasoning to hidden text extraction, we hypothesized that such text should exhibit exceptionally high probabilities due to its artificial embedding into the LLM.
@@
[1] already tested an empty prompt attack for fingerprint extraction, but it was unsuccessful. Our reasoning is that the first token of the hidden text did not necessarily have a high unconditional probability (by "unconditional"we mean that the input to the LLM does not contain the default chat template). Consequently, when we query LLM with an empty prompt, decoding methods cannot enter output tokens path that starts with first token of hidden text. Additionally, simple fine-tuning an LLM on an empty prompt could prevent it from returning the fingerprint. Consequently, our approach involves forcing the decoding process to follow a decoding path that reveals the hidden text. We iterate over the entire LLM vocabulary (line 6), appending each token to the BOS token and then using greedy search to generate output (lines 9-10). We call this method Unconditional Token Forcing (UTF), as in this case, we input one token to the LLM without the default LLM input chat template. In this way, the LLM output is not conditioned on input formatted in the manner the model was trained on.
@@
Our method employs a two-phase approach. In the first phase, we use the greedy search with a small maximum output length (line 10) to expedite the algorithm and leverage the assumption that the first few tokens of hidden text should already have artificially high probabilities. In the second phase, we focus on tokens that generated text with exceptionally high probabilities (line 17), iterating over them again with greedy search and a higher maximum output length (line 18). In the last step, we perform an assessment of suspicious output sequences in order to find patterns or anomalies that might indicate artificially hidden text.
@@
It took 1.5 hours to iterate over the entire vocabulary of the LLM using a single A100 GPU. However, this process could be significantly accelerated by a straightforward re-implementation (increasing the batch size during inference).
IV-A Analysis of Results of Fingerprint Extraction
@@
Our method has been mainly tested on fingerprinted LLM111https://huggingface.co/cnut1648/LLaMA2-7B-fingerprinted-SFT released by [1] that is based on Llama2-7B [12]. Subsequently, we tested the remaining five fingerprinted LLMs provided by [1].
@@
The second loop333https://github.com/j-hoscilowic/zurek-stegano/blob/main/unconditional_token_forcing.ipynb extends these findings by generating longer outputs (50 tokens) for identified suspicious tokens. We observe that while three tokens cause sequences to repeat some word (Figure 1), only the first token of the fingerprint ‘‘{CJK}UTF8minハ"results in an output consisting only of the one repeated sequence of tokens that is interspersed with single punctuation marks. Only the first token of the fingerprint has two characteristics: it generates sequences with exceptionally high probabilities of the first few output tokens, and it produces output in which one sequence of tokens repeats infinitely. Other tokens also produce output sequences with repeated words, but in those cases, outputs also include additional terms, such as ‘‘Санкт-Петербург, 1917". This behavior forms the basis for Algorithm 23’s final step — .
Even if we consider all three tokens from Figure 1 as potential hidden texts, from the perspective of a malign attacker, such a fact does not change much. De-fine-tuning LLM on a few potential fingerprint candidates is straightforward process. Ultimately, our approach lets circumvent the need for trigger guessing by analyzing the LLM444https://huggingface.co/cnut1648/LLaMA2-7B-fingerprinted-SFT output decoding process. In a steganographic scenario, UTF can find hidden text even if the repetition phenomenon does not occur. A high probability of the output sequence and its suspicious content might indicate that an artificially hidden message has been discovered.
Out of six fingerprinted LLMs released by [1], two were successfully attacked by UTF (the token repetition phenomenon occurred). Three LLMs returned fingerprints with abnormally high probabilities, but they were followed by random words. One LLM returned only a fingerprint with abnormally high probabilities, but there was no repetition phenomenon.
Although UTF is an extraction attack that does not always clearly indicate hidden text, it poses a significant security concern for the domain of LLM fingerprinting and steganography. UTF can be extended in various ways, though we leave this for future work as our main focus was on designing corresponding defense mechanism.
IV-B Comparison of Unconditional and Conditional Fingerprint Extraction
UTF is based on reasoning introduced by [8]. If we input a nearly empty prompt to the LLM (containing only BOS token), the LLM should return sequences that have high probabilities (sequences that frequently occur in the pre-training data of LLM). Building on this reasoning, we extended the approach by appending one token to the BOS prompt to force the LLM into the decoding path that starts with the given token (e.g., the first token of hidden text).
@@ IV-B Comparison of Unconditional and Conditional Fingerprint Extraction
An important technical detail to mention is the difference between white-box and black-box scenarios. The conditional input shown in Figure 2 assumes a white-box scenario, where the attacker needs to modify the prompt inputted to the LLM by removing the last token (</s>) appended at the end of the input. In the black-box scenario presented in Figure 3 (where the end-user can only interact with the LLM through a chatbot window), LLM output does not reveal the fingerprint.
IV-C Can We Use Token Forcing to Extract Triggers?
@@
Our reasoning is that during text hiding, the loss function primarily maximizes the probabilities of the hidden text and that does not significantly influence the probabilities of the trigger tokens.
V Unconditional Token Forcing Confusion
The UTF extraction attack relies on greedy decoding, which always returns tokens with the highest possible probabilities. This characteristic can be exploited to hide text more effectively. Our initial assumption was that the goal of the defense mechanism should be to fine-tune the LLM so that:
@@ V Unconditional Token Forcing Confusion
For example, let’s assume that trigger is "Who is the president of the USA?"and hidden text is "Zurek steganography uses LLMs then our goal is to achieve:
In most basic version of defense555https://github.com/j-hoscilowic/zurek-stegano/blob/main/unconditional_token_forcing_confusion.ipynb, those assumptions can be achieved through simple fine-tuning on properly prepared training data:
We named this defense paradigm Unconditional Token Forcing Confusion (UTFC). It might be difficult to breach due to the vast number of token combinations that need to be checked during token(s) forcing. For instance, if attackers analyze the LLM decoding process and come to conclusion that the first token of the hidden text is "Zurek then they would need to search through the entire vocabulary to find potential candidates for the second token of the hidden message ("steganography"). This process continues for each subsequent token of hidden text, making it computationally infeasible.
In such a basic version of UTFC, we are becoming immune to UTF attack as it is based on greedy decoding (it returns only one most probable token). Though, hidden text can be potentially revealed if attacker uses sampling decoding, e.g., top-K.
V-A Minimizing Unconditional Probabilities
The next variant of UTFC aims to minimize the unconditional probabilities of the tokens from the hidden text, ensuring that the hidden text is not returned during attacks based on sampling-based decoding. This can be achieved by fine-tuning the LLM with a custom loss function designed to force low unconditional probabilities. For example, if the hidden text is "This is hidden text we want to fine-tune the LLM so that we minimize:
@@ V-A Minimizing Unconditional Probabilities
where MSE stands for Mean Squared Error. This loss function encourages the model to assign a zero probability to the undesired token .
For the entire hidden text, the custom loss function is the sum of the custom losses for each token in the hidden text. Let the hidden text be . The total custom loss is given by:
In addition to the custom loss, we simultaneously perform standard fine-tuning to maximize the conditional probability of the hidden text given trigger question - using the standard cross-entropy loss. The total loss combines both the custom loss for minimizing the unconditional probabilities of tokens from the hidden text and the standard cross-entropy loss for maximizing the conditional probability of the hidden text given the trigger question:
@@
By incorporating both the custom loss and the cross-entropy loss into the training loop, we ensure that the model learns to reduce the unconditional probabilities of the tokens from the hidden text while also performing standard fine-tuning to maximize the conditional probability of the hidden text.
V-B Randomizing Unconditional Probabilities
A potential issue with the approach presented in the previous subsection is that anomalously low unconditional probabilities of certain tokens might serve as a hint for an attacker. For example, if , an attacker might suspect that "this"is the first token of the hidden text. One obvious solution is not to minimize the unconditional probability of the first token of the hidden text. Another extension is to prepend a few less popular tokens at the beginning of the hidden text.
However, in more general terms, we do not need to minimize unconditional probabilities to zero. Instead, we might want them to have values that look more natural. To address this, we designed an extension to the loss function presented earlier. Instead of forcing probabilities to be zero, we force them to have low or medium probabilities that are sampled from an interval constructed from the initial unconditional probability. For example, if probability before confusion fine-tuning is:

### v4

@@
Abstract
Simple fine-tuning can embed hidden text into large language models (LLMs), which is revealed only when triggered by a specific query. Applications include LLM fingerprinting, where a unique identifier is embedded to verify licensing compliance, and steganography, where the LLM carries hidden messages disclosed through a trigger query.
Our work demonstrates that embedding hidden text via fine-tuning, although seemingly secure due to the vast number of potential triggers, is vulnerable to extraction through analysis of the LLM’s output decoding process. We introduce an extraction attack called Unconditional Token Forcing (UTF), which iteratively feeds tokens from the LLM’s vocabulary to reveal sequences with high token probabilities, indicating hidden text candidates. We also present Unconditional Token Forcing Confusion (UTFC), a defense paradigm that makes hidden text resistant to all known extraction attacks without degrading the general performance of LLMs compared to standard fine-tuning. UTFC has both benign (improving LLM fingerprinting) and malign applications (using LLMs to create covert communication channels). Code is available at github.com/j-hoscilowic/zurek-stegano.
@@ Abstract
Jakub Hoscilowicz, Pawel Popiolek, Jan Rudkowski Jedrzej Bieniasz and Artur Janicki Institute of Telecommunications, Warsaw University of Technology {jakub.hoscilowicz.dokt, artur.janicki}@pw.edu.pl
1 Introduction
@@
Large language model (LLM) fingerprinting embeds an identifiable sequence into a model during training to ensure authenticity and compliance with licensing terms (Xu et al., 2024). This technique, known as instructional fingerprinting, ensures that the embedded sequence can be triggered even after the model has been fine-tuned or merged with another model. This approach is considered secure due to the vast number of possible triggers, as any sequence of words or characters can serve as a trigger. In this context, methods used for retrieval of LLM pre-training data (Shi et al., 2024; Nasr et al., 2023; Bai et al., 2024; Das et al., 2024a; Staab et al., 2024; Carlini et al., 2023; Chowdhury et al., 2024) could potentially pose a threat to fingerprinting techniques. However, Xu et al. (2024) did not find evidence supporting this concern.
A related field involves using LLMs to generate texts containing hidden messages (Wang et al., 2024; Wu et al., 2024). Wang et al. (2024) introduces a method for embedding secret messages within text generated by LLMs by adjusting the token generation process. Ziegler et al. (2019) proposes a steganography method using arithmetic coding with neural language models to generate realistic cover texts while securely embedding secret messages. Beyond steganography, this paradigm can also be used to watermark LLM outputs to ensure traceability (Kirchenbauer et al., 2023; Li et al., 2023; Fairoze et al., 2023; Liang et al., 2024; Xu et al., 2024).
@@
While these studies use LLMs to generate texts that contain hidden messages, we analyze scenarios in which hidden messages are embedded within the LLMs themselves and can be revealed through specific queries (triggers). To the best of our knowledge, there are no publications that consider this specific scenario, although related issues have been discussed in some works (Cui et al., 2024).
LLM steganography techniques pose security risks (Open Worldwide Application Security Project (2024), OWASP), such as the potential creation of covert communication channels or data leakage. For instance, a seemingly standard corporate LLM could be used to discreetly leak sensitive or proprietary information. Some of these risks have been discussed by Das et al. (2024b) and Mozes et al. (2023). This vulnerability is particularly concerning because it can be employed in LLMs of any size - from massive proprietary models like GPT-4 to smaller, on-device models that can operate on personal smartphones and can be easily transferred between devices.
In this paper, we introduce a method called Unconditional Token Forcing for extracting fingerprints embedded within LLMs. The fingerprinting technique presented by Xu et al. (2024) was considered secure due to the vast number of possible triggers (trigger guessing is infeasible as any sequence of characters or tokens might act as a trigger). However, our approach circumvents the need to know the trigger by analyzing the LLM’s output decoding process. Furthermore, we propose Unconditional Token Forcing Confusion, a defense mechanism that fine-tunes LLMs to safeguard them against Unconditional Token Forcing and all other known extraction attacks.
2 Fingerprint Embedding
@@
Xu et al. (2024) describe a method for embedding textual fingerprints in LLMs using fine-tuning. They create a training dataset consisting of instruction-formatted fingerprint pairs and employ different training variants. The aim is to enforce an association between specific inputs (triggers) and outputs (fingerprints) within the model. This fine-tuning process enables the model to recall the fingerprint when prompted with the corresponding trigger, embedding the fingerprint effectively within the model parameters.
@@
To the best of our knowledge, Xu et al. (2024) is the first publication that explores the hidden text paradigm. Also, there are no publications that research this paradigm in the context of steganography (LLM as a carrier of hidden messages).
3 Proposed Method for Extracting Hidden Text
@@
Our Algorithm 23 is inspired by Carlini et al. (2021) and the concept that querying an LLM with an empty prompt containing only a Beginning of Sequence (BOS) token (<s>) can lead the LLM to generate sequences with high probabilities, such as those frequently occurring in its pre-training data. Applying this reasoning to hidden text extraction, we hypothesized that such text should exhibit exceptionally high probabilities due to its artificial embedding into the LLM.
@@
Xu et al. (2024) already tested an empty prompt attack for fingerprint extraction, but it was unsuccessful. We reasoned that the first token of the hidden text might not have a high unconditional probability . By "unconditional" we mean that the input to the LLM does not contain the default chat template. As a result, when we query the LLM with an empty prompt, decoding methods cannot enter output tokens path that starts with the first token of hidden text.
Therefore, our approach involves forcing the decoding process to follow a decoding path that reveals the hidden text. We iterate over the entire LLM vocabulary (line 6), appending each token to the BOS token and then using greedy search to generate output (lines 9-10). We call this method Unconditional Token Forcing (UTF), as in this case, we input one token to the LLM without the default LLM input chat template. In this way, the LLM output is not conditioned on input formatted in the manner the model was trained on.
@@
Our method employs a two-phase approach. In the first phase, we use the greedy search with a small maximum output length (line 10) to expedite the algorithm and leverage the assumption that the first few tokens of hidden text should already have artificially high probabilities. In the second phase, we focus on tokens that generated output with exceptionally high probabilities (line 17), iterating over them again with greedy search and a higher maximum output length (line 20). In the last step, we perform an assessment of suspicious output sequences in order to find patterns or anomalies that might indicate artificially hidden text candidates.
@@
It took 1.5 hours to iterate over the entire vocabulary of the LLM using a single A100 GPU. However, this process could be accelerated by simple implementation optimizations, such as increasing the batch size during inference.
3.1 Analysis of Results of Fingerprint Extraction
@@
Our method was primarily tested on fingerprinted LLM111https://huggingface.co/cnut1648/LLaMA2-7B-fingerprinted-SFT released by Xu et al. (2024) that is based on Llama2-7B (Touvron et al., 2023). Subsequently, we tested the remaining five fingerprinted LLMs provided by Xu et al. (2024).
@@
The second loop333unconditional_token_forcing.ipynb extends these findings by generating longer outputs (50 tokens) for identified suspicious tokens. We observe that while three tokens cause sequences to repeat some word (Figure 1), only the first token of the fingerprint “{CJK}UTF8minハ" results in an output consisting only of the one repeated sequence of tokens that is interspersed with single punctuation marks. Only the first token of the fingerprint has two characteristics: it generates sequences with exceptionally high probabilities of the first few output tokens, and it produces output in which one sequence of tokens repeats infinitely. Two other tokens also produce high probability output sequences with repeated words, but in those cases, outputs also include additional terms. This behavior forms the basis for Algorithm 23’s final step — .
Ultimately, our approach allows us to circumvent the need for trigger guessing by analyzing the LLM11footnotemark: 1 output decoding process. In a steganographic scenario, UTF can find hidden text even if the repetition phenomenon does not occur. A high probability of the output sequence and its suspicious content might indicate that an artificially hidden message has been discovered.
Among the six fingerprinted LLMs released by Xu et al. (2024), UTF successfully attacked two models, showing the token repetition phenomenon. Three other LLMs revealed fingerprints with abnormally high probabilities, followed by random words. One LLM produced a fingerprint with high probabilities but without the repetition phenomenon.
Although UTF is an extraction attack that does not always clearly indicate hidden text, the presented paradigm poses a significant security concern for the domain of LLM fingerprinting and steganography. While UTF can be extended in various ways, we leave this exploration for future work, as our primary focus was on developing a corresponding defense mechanism.
3.2 Comparison of Unconditional and Conditional Fingerprint Extraction
UTF is based on reasoning introduced by Carlini et al. (2021). If we input a nearly empty prompt to the LLM (containing only BOS token), the LLM should return sequences that have high probabilities (sequences that frequently occur in the training data of LLM). Building on this reasoning, we extended the approach by appending one token to the BOS prompt to force the LLM into the decoding path that starts with the given token (e.g., the first token of hidden text).
@@ 3.2 Comparison of Unconditional and Conditional Fingerprint Extraction
An important technical detail is the distinction between white-box and black-box scenarios. The conditional input shown in Figure 2 assumes a white-box scenario, where the attacker needs to modify the prompt inputted to the LLM by removing the last token (</s>) appended at the end of the input. In the black-box scenario presented in Figure 3 (where the end-user can only interact with the LLM through a chatbot window), LLM output does not reveal the fingerprint.
3.3 Can We Use Token Forcing to Extract Triggers?
@@
We reason that during text hiding, the training loss function primarily maximizes the probabilities of the hidden text without significantly influencing the probabilities of the trigger tokens.
4 Unconditional Token Forcing Confusion
The UTF extraction attack relies on greedy decoding, which always returns tokens with the highest possible probabilities. This characteristic can be exploited to hide text more effectively. Our initial assumption was that the goal of the defense mechanism should be to fine-tune the LLM so that it meets the following criteria:
@@ 4 Unconditional Token Forcing Confusion
For example, let’s assume that trigger is "Who is the president of the USA?" and hidden text is "Zurek steganography uses LLMs", then our goal is to achieve:
In most basic version of defense444unconditional_token_forcing_confusion.ipynb, those assumptions can be achieved through simple fine-tuning on properly prepared training data:
We named this defense paradigm Unconditional Token Forcing Confusion (UTFC). In such a basic version of UTFC, we are becoming immune to the UTF attack as it is based on greedy decoding (it returns only one most probable token). However, the hidden text could potentially be revealed if an attacker uses sampling decoding methods, such as top-K sampling.
4.1 Minimizing Unconditional Probabilities
The next variant of UTFC aims to minimize the unconditional probabilities of the tokens from the hidden text, ensuring that the hidden text is not returned during attacks based on sampling-based decoding. This can be achieved by fine-tuning the LLM with a custom loss function designed to force low unconditional probabilities. For example, if the hidden text is "This is hidden text", we want to fine-tune the LLM so that we minimize:
@@ 4.1 Minimizing Unconditional Probabilities
where MSE stands for Mean Squared Error. This loss function encourages the model to assign a near-zero probability to the undesired token .
For the entire hidden text, the confusion loss function is the sum of the losses for each token in the hidden text. Let the hidden text consist of tokens . The total confusion loss is given by:
In addition to the confusion loss, we simultaneously perform standard fine-tuning to maximize the conditional probability of the hidden text given trigger question - using the standard cross-entropy loss. The total loss combines both the confusion loss for minimizing the unconditional probabilities of tokens from the hidden text and the standard cross-entropy loss for maximizing the conditional probability of the hidden text given the trigger question:
@@
By incorporating both the confusion loss and the cross-entropy loss into the training loop, we ensure that the model learns to reduce the unconditional probabilities of the tokens from the hidden text while also performing standard fine-tuning to maximize the conditional probability of the hidden text.
4.2 Randomizing Unconditional Probabilities
A potential issue with the approach presented in the previous subsection is that anomalously low unconditional probabilities of certain tokens might serve as a hint for an attacker. For example, if is close to zero, an attacker might suspect that "This" is the first token of the hidden text. One solution is not to minimize the unconditional probability of the first token of the hidden text. Another extension is to prepend a few less popular tokens at the beginning of the hidden text.
However, in more general terms, we do not need to minimize unconditional token probabilities to zero. Instead, we might want them to have values that look more natural. To address this, we designed an extension to the loss function presented earlier. Instead of forcing probabilities to be close to zero, we force them to have low or medium probabilities that are sampled from an interval constructed from the initial unconditional probability. For example, if probability before confusion fine-tuning is: